Journalist Report – January 22nd

Journalist Report:


[Sol 10]

[Apollo 13] [Eng]

With a quiet morning, the Sun 10 began, breakfast of sugary cereal and orange flavored drinks. Then some Martians disguised specialists in terrestrial habitats visited us to take charge of matters related to failed systems. It seems that from today our vital fluid will be safer and more reliable. Showers with hot water will be possible now.

Four of our bravest crew went out in search of the conquest of Mars aboard two rovers. An hour later the radios announced their return, in the habitat we worried. A rover’s battery failed and the cold froze the astronauts’ hands.

After the change of vehicle and warm up better they headed towards Copernicus Highway. Upon his arrival, amazement shinned their eyes. Wherever we go, we are the first to be there. It is the thought that unite our four crew members after traveling to the adventure to know and carry out their investigations. Result: a planetary wonder.

What did they find? A prehistoric land that surrounded in the four cardinal points like being in the middle of a red sea and that as in the movies that we watched in the big screen a dinosaur could appear in the middle of the canyons. Adrenaline and pure emotion.

Meanwhile the habitat consisted of one day of cleaning and inventory of surplus food for the remaining days on Mars. After climbing up and down the wooden stairs we have become accustomed to the feeling of living in a tree house as we did in our childhood.

But the adventure of our explorers did not end there until much later. On the way back another of the rovers consumed his battery until almost totally static. Fortunately we have a prepared and alert crew that managed to reach our home. Although the final result was positive, the road took more than two hours and consumed its energy and heat.

To get home they used brute force and ingenuity between tying ropes, cables, almost flags and pushing flags. Within five minutes of entering the habitat, they were rescued by the Station Director and finally returned safely.

We received them with scarves, sweaters, blankets and hot chocolate with marshmallows. History and enthusiasm have distracted us a lot of time tonight, but luckily we are all together and excited to continue exploring our red earth. There is no doubt that what makes us human, friendship, companionship and love is what will always get us out of trouble, here, there and on any planet, with faith.

LATAM II will continue to inform

Tania Robles, MDRS Crew 187 Journalist

Sol Summary – January 22nd

Crew 187 Sol 10 Summary Report 22JAN2018

Sol 10

Summary Title: Bring it on home to me

Author’s name: Cynthia Fuertes Panizo

Mission Status: All nominal

Sol Activity Summary:

Houston… we have a problem!

Today we decided to wake up in order to rest for the EVA of today. Every Martian woke up in different hours. The majority of them were taking a breakfast at 08:15, but there is always someone who falls asleep and gets up late, at 9:30, and woke up only by the noise that made two earthlings who visited our home on Mars to fix the water heater and water filter. Due to this, 130 gallons of water were used today… Earth, we need more water to survive on Mars!

In the morning, Oscar was working in the programing phase of his project, and in the afternoon his rover was tested inside the tunnels. It’s magic how the rover can move through rocks, sands and can climb the tunnel’s walls. David and I were working in our projects in the Greenhab. Moreover, Atila is trying to fix the Science Dom’s door. At 10:35, every Martian were in the upper deck to help each other to prepare for the EVA of today. At 12:00, David, Atila, Luis and Danton were in the airlock ready for their adventure. However, at 12:47 Hab copied a radio signal from the Martian explorers. Due to the fact that Opportunity charger was getting down fast; in 18 minutes the battery got down more than 10% arriving until 64%. And prioritizing the safety of the crew’s members, it was decided to come back to the Hab. The decision taking was that the explorers got inside the Hab, ate a lunch, warmed up, and started again the EVA. Oscar cooked very quickly eggs with cheese for the Martian explorers of the day. For the second excursion, Deimos and two 350 ATV were being used. In the afternoon, Tania, Oscar and I cleaned the Kitchen and did a food inventory. We found some interesting food that we can eat in the following days. At 16:45 Hab didn’t know anything about the Martian explorers and we started to worry about it. At 16:56, Hab copied the first radio signal and we realized that every Martian was ok but something happened. At 14:50, the Martian explorers arrived to Copernicus Hwy. The Martians described the view as the best landscape that they have ever seen in their lives. At 13:50, the Martian explorers were ready to come back to home. At 16:45 Deimos stoped and the Martian worked together using jumping cables to tow Deimos, that’s why they arrived to Hab at 18:20. At 17:45, Shannon found the Martian explorers and helped them to solve the problem.

The day was a really crazy day… but come on! We are on Mars! We are Martians! And the important thing is that we knew how to solve it!

Ad Astra,

Cynthia Fuertes Panizo

Commander Crew 187 – MDRS

Look Ahead Plan:

We hope to have another adventure for tomorrow… We are going to Lith Canyon! The lucky Martians are Oscar, Tania, Cynthia and Danton.

Anomalies in work:

It’s the Sol Summary.

Crew Physical Status: Healthy

Reports to be filed: Sol Summary, GeenHab Report, Operations Report, EVA Request, Science Report, Astronomy Report and Journalist Report.

Support Requested:

Please send us more jumping cables and a refill of water.

Greenhab Report – January 22nd

GreenHab Report

Hernán David Mateus Jiménez


Environmental control:


Ambient with door opening

Shade cloth on

Working Hour: 20:01
Inside temp at working hour: 18° C
Outside temp during working hours: -4° C
Inside temperature high: 31° C
Inside temperature low: 14° C
Inside humidity: 80 %RH

Inside humidity high: 86 %RH
Inside humidity low: 16 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: The carrots and onions started to grow

Daily water usage for crops: 8 gallons

Time(s) of watering for crops: 19:15

Changes to research plants: none

Aquaponics: Atila is assembling it. He is waiting for the amount of water to fill the tank, this is necessary to do the test.

Narrative: None

EVA Report – January 22nd

This is the report for today’s EVA

Purpose of EVA: Primary: Long-range EVA protocol testing, Sample collection for possible Halophile identification, identification of relevant sampling sites, heating vest testing.

Secondary: Outreach and Educational graphic material collection. Relevant sample collection.

Participants: Atila, Danton, Luis, and David.

Coordinates: 517100 E, 4256400 N

Narrative: Today was the first long range EVA undertaken by the crew, and it was an eventful one, fortunately with a very happy ending. It is important to say that the temperature was a challenging element for the crew. The EVA started at 12:00 with the egress of the crew of four of the airlock. By pairs they boarded the two rovers and departed north by Cow Dung road, however, by the rapid discharge of the opportunity rover, the crew returned to the hab to rest, eat something, and take two ATVs instead of Opportunity. The EVA was resumed at 13:30 with a new egress of the crew, this time boarding Deimos and the 1 and 2 350 ATVs. The crew departed north until the north of Copernicus Hw, where they left the vehicles and started walking at 14:50. The crew walked about 15 minutes to a valley(38° 27.448N 110° 48.031W) and stayed there for about 5 minutes, where they took the route back, taking relevant samples in the way, as well as aerial footage(38° 27.0021N 110°48.565W). There were a lot of hervibore prints as well. While on Brahe Hwy the battery meter of Deimos suddenly started dropping fast, taking about 5 minutes to go from 71% to 26% charge, as reported by the Engineering officer. The crew decided to switch positions to reduce the weight, and kept going, however the charge kept dropping. The crew decided to walk along the vehicle, and afterwards to push it. The crew followed this procedure for 50 minutes. Afterwards they took the jumping cables (for battery charging) and pulled the rover with one of the ATVs until they were in sight of the crew in the hab. Unfortunately, the cables broke, and it was impossible to pull the rover further. At that point the crew was already 45 minutes above the closing time of the EVA window, and director Shannon went out to help them. The crew broke simulation, towed and pushed the rover up to the habitat and made an ingress to the hab, out of sim.

They were greeted with hot cocoa and blankets by the rest of the crew, who were happy to see them again.

Operations Report – January 21st

Operations Report

Subject Line: Crew 187 Operations Report 21JAN2018

Crew 187 Operations Report 21012018

SOL: 09

Name of person filing report: Luis Jose Diaz Lopez

Non-nominal systems: Water heater.

Notes on non-nominal systems: Generator system limping along with a now-nominal routine.

Generator (hours run): 15h 10min

Generator turned off, charging battery at 10h:58m

Generator turned on at 19h:00m

Solar— SOC

@ 08:42: 65%

@ 18:11: 69%

Diesel: It can’t be measure. Explained in the Report 19JAN2018

Propane: 95%

Ethanol Free Gasoline (5 Gallon containers for ATV): 4.75 Gallons

Water (trailer): 0 Gallons

Water (static): 230 Gallons

Trailer to Static Pump used: Yes

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 01300547 Gallons

Toilet tank emptied: No

ATVs Used: None

Oil Added? No

ATV Fuel Used: – Gallons

# Hours the ATVs were used today: –

Notes on ATVs: ATVs were nominal.

Deimos rover used: No

Hours: 101.8 h0

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

Sojourner rover used: Assigned to director only.

Hours: Director discretional hours

Beginning charge:

Ending charge: –

Currently charging: Maybe

Spirit rover used: Yes

Hours: 12.7h

Beginning charge: 100%

Ending charge: –

Currently charging: No

Opportunity rover used: Yes

Hours: 5.6h

Beginning charge: 100%

Ending charge:

Currently charging: Yes

Curiosity rover used: No

Hours: 39h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

HabCar used and why, where? No

General notes and comments: The power generator was running for several hours because there were many clouds in the morning that did not allow the solar panels to charge the batteries. Water heater was disconnected because it’s not working.

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal. Today we got a refill of 125 gallons of water from the water (trailer).

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: None

Sol Summary – January 21st

Crew 187 Sol 9 Summary Report 21JAN2018

Sol 9

Summary Title: Zombie

Author’s name: Cynthia Fuertes Panizo

Mission Status: All nominal

Sol Activity Summary:

Is there someone in Earth?

This day we decided not to do the exercises and start the day late, so at 8:15 we ate breakfast and every Martian started to work in their project, our alarm was delayed an hour. The tiredness begins to be noticed because the naps are more common between meals. At 11:00 our Martian chef, Danton, started to cooked a delicious food: Chicken with cheese, pasta and vegetables. We ate together before the expedition began. At 11:30, Tania, David, Luis and Oscar started the donning time. They took Opportunity and Spirit for this adventure… It’s supposed that they were going to work in El Dorado Canyon, but it was a confusion with the location so they visited a near canyon. However, they took some samples of the halophiles needed. The path that they follow was bent. The canyon that the Martians explorers visited combines a wide variety of landscapes, from rocks up to 5 meters high, which route land for an approximate of 500 meters between ascents and descents and with different types of soil and frozen water remains.

By 15:00, our chef Danton started to cooked a second lunch for our Martian explorers, a delicious soup and scrambled eggs with chili. They arrived at 15:20 to our home. And 25 minutes later we started to eat together… again!

In the afternoon, we got a refill of 125 gallons from the water (trailer). Now we have water for the rest of our lives in Mars, or at least, we hope so.

David is replanting 12 quinua’s plants on the lysimeter for his project. Meanwhile, I get an advanced programing my augmented reality mobile application. And, as always, Tania is documenting everything for her project.

For dinner, Atila is trying to cook a surprise Martian dinner…. Everybody is waiting to taste it!

Ad Astra,

Cynthia Fuertes Panizo

Commander Crew 187 – MDRS

Look Ahead Plan:

Another EVA is going to take place. The lucky ones are Luis, Atila, David and Danton. They are going to Copernicus Hwy.

Anomalies in work:


Weather: Temperature: -1°C, Fair weather, Winds speed: NNE 8 KPH, Humidity 43%, Barometer 30.22 in

Crew Physical Status: Healthy

Reports to be filed: Sol Summary, GeenHab Report, Operations Report, EVA Request, EVA Report and HSO Report.

Support Requested:

Please sent us distillated water.

GreenHab Report – January 21st

GreenHab Report

Hernán David Mateus Jiménez


Environmental control:

Shade cloth on

Working Hour: 18:20
Inside temp at working hour: 19° C
Outside temp during working hours: -1° C
Inside temperature high: 24° C
Inside temperature low: 15° C
Inside humidity: 48 %RH

Inside humidity high: 48 %RH
Inside humidity low: 16 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: None

Daily water usage for crops: 8 gallons

Time(s) of watering for crops: 18:30

Changes to research plants: none

Aquaponics: Atila is assembling it. He is waiting for the amount of water to fill the tank, this is necessary to do the test.

Narrative: The lysimeter is taking data. Today 12 quinoa plants were replanted in the lysimeter over Martian soil. Atila and I have made some experiments to discovers what is a good proportion between soil and water to get the best texture, we have figured out that you need 500 mL of water for 2800 g of soil.

EVA Report – January 21st

Purpose of EVA: Primary: Mid-range testing of the Cóndor Space Suit Simulator, Mid-range EVA protocol testing, Sample collection for possible Halophile identification, further exploration of El Dorado Canyon.

Secondary: Outreach and Educational graphic material collection. Relevant sample collection.

Participants: Oscar, Luis, David, and Tania.

Narrative: Today marked the last of the mid-range EVAs programmed, leading to longer and even more interesting exploration. The weather was favorable, but windy. The crew donned the suits and exited the airlock at 12:30, the four crewmembers then boarded the Rovers Spirit and Opportunity and followed the path planned, towards El Dorado canyon. After approximately 25 minutes of traverse, and after passing Robert’s Rock Garden, the Rovers were parked on the side of the road, and the expedition started walking with Southeast course, towards the entrance of the canyon, upon finishing the rocky plains, two crewmembers had to return to the Rover, given that the aerial drone was left behind. Upon reuniting, the crew resumed the traverse until the entrance of the canyon. More herbivore and cougar prints were found in the area, but not as fresh as on the previous EVA to the canyon. The crew entered the canyon and followed the dry riverbed. Along this path, water samples were taken, aerial footage with the drone, as well as traditional photography. The crew followed the path until the eastern exit of the canyon, heading south to meet Cow Dung road again. However, the sampling sites that were found during the firs EVA to the area were not found, since they probably belong to a subsidiary structure, south of El Dorado canyon. It is suggested that further detailed mapping of this area is undertaken, to identify all the morphologies present. While returning through Cow Dung road, heading north, a crewmember’s cell phone (used as a compass and camera) was lost, so a search formation was done, and the crew headed south, being able to find the device, and resuming the walk back north. The walk lasted about 25 minutes, reuniting with the Rovers and heading back to the habitat. The EVA ended at 15:20 with the ingress of the crew from the airlock to the habitat.

The aerial footage is being processed, as well as the performance of the Cóndor Space Suit Simulator, which will be subjected to detailed inspection tomorrow.

The only incident related with the EVA was the lost and recovery of the cell phone.

Science Report – January 20th

Good evening Earthlings,

A summary of the crewmembers research follows:



This drone will map the soil of the MDRS during two weeks which will help to take images at 40 meters of height to be later analyzed by a digital processing software which will help us to better understand the characteristics of the Mars surface as well the automatic pilot of drone will help to astronauts recollecting information in difficult areas to explore. After that Using digital image processing algorithms, we will determine the characteristics of the surface of the MDRS using Matlab to analyze the images taken by the drone.



This device will monitor with the E.C.G sensor as well as some important aspects like the pressure before each EVA of Crew 187 LATAM-II. This device will be able to measure the internal temperature of the astronaut as well as the humidity inside the suit in addition ECG module moreover the body position, galvanic response skin that will transmit the data to the user interface to PC and will get the support to Crew 187.



Analyze the dynamics of different cognitive abilities and its relationship with fatigue levels during the mission, in astronauts and to compare them against a matched control group of external participants not related to the Analogue Simulation. Fourteen adults will be part of the study. A control/cases design will be employed. 7 subjects that are part of the astronauts group of Crew 187 LATAM-II and seven persons paired to every participant from the cases group will be used as the control group. The groups will be paired by the sex, age, lateral dominance and level of studies. The subjects from the support group and the control paired will be chosen voluntarily. The study will be divided in three stages: “pre-mission”, “during-mission” and “post-mission “In each stage, fourteen participants will solve assigned COGNIFIT and Multidimensional Fatigue Test.



Analyze the dynamics of cooperation and reciprocity between the Analogue Simulation Crew 187 LATMA-II members 2 women and 4 men between 23 and 30 years old will be part of the study. After that the information of the cooperative behavior between crew members during the analogue simulation to Mars, will be apply a Collective-Risk Social Dilemma in which six astronauts will be players and one coordinator. This task will be applied five times for two weeks.


Our search for the lovers of extreme just began. Halophiles are organisms that can flourish on high salinity environments, and MDRS sorroundings are the perfect place to haunt them. We’ve made a Halophile Sampling Method Protocol for tomorrow’s EVA to perform a regular sampling on El Dorado Canyon and it will be useful for our next explorations on Copernicus Highway.

The aquaponics isn’t running yet, but it’s almost completely build up. New Moringa seeds are coming pretty soon. We are waiting to do some equipment tests and after that, only one week will be necessary to start running it.

Today Atila’s project related to quinoa and kiwicha just began, pictures showing the planted pots with the seeds on the Mars analog soil are attached. We are looking forward to seeing those babies germinate. Previous germination test suggests that we are expecting seedlings for Sol 12. We will let you know

Researcher: Luis Jose Antonio Diaz Lopez

It was checked that the control modules: DS18B20 + RTC DS3132 + SD CARD are working correctly. The programming of the microcontroller is ready to perform the measurement of the two temperature sensors to obtain the values inside and outside the suit used during the EVA’s. The data sensors are transmitted to a cell phone via a bluetooth module and also store them in a .csv file inside a microsd card for further processing. Tomorrow (SOL 9) it will be tested in a EVA.

Researcher: Tania Robles

Science communication and documentary to space projects of young scientist and professionals in Latin America.

Progress: The documentary process on the participation of young people in space science and technology in Latin America has advanced. According to the observation of the behavior and participation of the crew members, videos and notes about each member have been taken.

Today, the first personal interview was held with the GreenHab Officer, where personal and professional issues related to their projects, decisions and future in the space area in their country and in the international community have been addressed.

In the following days, the missing members will be interviewed and shots will be taken in each one of the work and development environments of the crew members.

Researcher: Oscar Ojeda

Progress: There are two projects in course, the testing of the Cóndor Space Suit Simulator, and the Testing of the PXCM based wheel for a planetary rover. All systems had to be transported completely unassembled, so the first task was related with the assembly and testing of them. At this point the Rover is completely assembled and the software will be tested in the next two Sols. The Space Suit Simulator has already been assembled and tested on short range and mid-range EVAs, the systems of the suit are all nominal, and further testing on mid and long range EVAs will be undertaken in the last 7 Sols.

Researcher: David Mateus

The lysimeter has been assembled and calibrated, in this moment it is tested with Martian soil and without plants. tomorrow It is expected to replant quinoa over the soil that it is tested. The idea is to measure evapotranspiration over this soil, in order to make a list of recommendations to improve the conditions on the greenhab, with the objective of reducing this phenomenon.

Researcher: Cynthia Yacel Fuertes Panizo

Mobile application as help agent in MDRS


The purpose of the application is to benefit the majority of the crew members, that’s why I analyzed the market share of mobile operating systems worldwide, a report published quarterly by Gartner, a leading company in market research. worldwide in the technology area. Therefore, the app will be developing to Android.

In order to choose the software of the tools that will be necessary like game engine, IDE (Integrated Development Environment) and augmented reality SDK, the following methodology will be use: First, list the different kind of software of each tool. Second, get the list of the top software. The third step is composed by two parts quantitative and qualitative analysis. Finally, the software of the tools is selected. In this way, Unity (game enginee), Monodevelop (IDE) and Vuforia (augmented reality SDK) are chosen (1).

One of the outputs of this project is an app for the crew astronomers. The target of this app will be the Hand Control of the Musk Observatory, and as it’s appreciate in the picture attached, the app already recognizes the real Hand control.

(1) C. Fuertes Panizo, “Aplicativo móvil de realidad aumentada para mejorar el proceso de enseñanza – aprendizaje”, 2017.